A particular fruit's weights are normally distributed, with a mean of 515 grams and a standard deviation of 14 grams.
If you pick 11 fruit at random, what is the probability that their mean weight will be between 501 grams and 502 grams (Give answer to 4 decimal places.)

Answer by Boreal(15235)

z=( x bar-mean)/sigma/sqrt(n)
=(501-515)/14/sqrt(11)=-3.317
and (502-515)14/sqrt(11)=-3.080
probability is 0.00058 or 0.0006.
check with (2ndVARS2 normalcdf(501,502,515,14/(sqrt(11)) ENTER